a particle is moving in a straight line with initial velocity u and retardation av, where v is the velocity at any time t

Grade:12th Pass

1 Answers

Aravind Bommera
11 years ago

Uniform Motion

 


In motion, we have studied about what are the parameters needed for a person or a body to move.
But, if a person is moving or any body is moving, how much will it cover in a given time? If we consider an instant of time, Is it moving uniformly or taking rest and then moving ? What makes a body move uniformly ?
To know the answers of all these questions, we will study the concept of Uniform Motion. The change in position of an object with respect to time is called as motion, which is measured with respect to a reference frame. Velocity, acceleration, displacement and time are the basic parameters which describe motion.

There are basically two types of motion :
  1. Uniform motion
  2. Non-uniform motion    
 

What is Uniform Motion?


A motion which covers equal distance in equal interval of time is called a Uniform motion.
For the body to be in the uniform motion, it must be moving in the straight line path.
Uniform motion

The above graph shows, in every one second, there is a displacement of 10m.
The body going with constant increase in velocity in equal interval of time is also the uniform motion.
The slope of displacement-time graph for uniform motion is constant and gives constant velocity.
Slope = $\frac{dY}{dX}$ = $\frac{displacement}{time}$ 

Velocity (V) = $\frac{d}{t}$, is constant


Uniform Acceleration Graph

Uniform acceleration

The V-t graph for uniform motion gives the constant acceleration.
The slope of uniform motion of V-t graph gives acceleration.
Slope = $\frac{dY}{dX}$ = $\frac{dv}{dt}$, gives acceleration. The SI unit is m/s2.

Some examples of uniformly accelerated motion:
  • The motion of a free falling body.
  • The motion of a bicycle going down the slope of a road when the rider is not pedaling and wind resistance is negligible.
  • The motion of a ball rolling down an inclined plane.
  • The motion of the Pendulum Clock

Equations of Motion


(a) V = U + at

(b) V
2 – U2 = 2aS

(c) S = U t + $\frac{1}{2}$ at
2

where V = Final Velocity
U = Initial velocity
a = Acceleration and
S = Displacement